# VideoChat-R1: Enhancing Spatio-Temporal Perception via Reinforcement Fine-Tuning

### v3

@@ 3.1 Preliminary of Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) [25] is a variant of Proximal Policy Optimization (PPO) [24] in reinforcement learning. By comparing groups of candidates responses directly, GRPO eliminates dependency on a critic model and significantly lowers training resources. Given an input question $q$, GRPO first generates $G$ distinct candidate responses $o=\{o_{1},\dots,o_{G}\}$ through policy sampling. The MLLM serves as the reward function to get the corresponding scores $\{r_{1},\dots,r_{G}\}$. GRPO computes their mean and standard deviation for normalization and determines the quality of these responses:
@@
[…]
$\max_{\pi_{\theta}}\mathbb{E}_{o\sim\pi_{\theta_{\mathrm{old}}}(p)}\Big{[}\sum_{i=1}^{G}\frac{\pi_{\theta}(o_{i})}{\pi_{\theta_{\mathrm{old}}}(o_{i})}\cdot A_{i}-\beta\,\mathrm{D}_{\mathrm{KL}}\Big{(}\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\Big{)}\Big{]}\text{,}$ (2)
where $\beta$ is a regularization coefficient, preventing excessive deviation from the reference policy during optimization.
@@ 3.2 Spatio-Temporal Rewards of Video MLLM in GRPO
$R_{\mathrm{format}}=\begin{cases}0,&\text{if output matches format},\\ 1,&\text{if output doesn't match format}.\end{cases}$ (3)
@@ 3.3 Enhance Spatio-Temporal Perception of Video MLLM through GRPO
Training data.
For the temporal grounding task, we use the training set of Charade - STA [8] (5,338 samples) for training. For the object tracking task, training is conducted on the GoT - 10k [11] dataset, which has 9,335 samples. For the QA and grounding QA tasks, the validation set of NExTGQA  [34] (3,358 samples) is used for training. For video captioning, FIBER-1k [35] (1,000 samples) is adopted for training. For video quality assessment, we use the quality assessment task from VidTAB [17] under the 100-shot setting, with 200 samples for training. Finally, for the training of VideoChat-R1, we perform joint training on three spatio-temporal perception-related tasks: temporal grounding, object tracking, and grounding QA. In total, 18,031 samples are used for training.
@@ 4.1 Evaluation of VideoChat-R1
Method Charades-STA ActivityNet NExTGQA GoT VideoMME MVBench Peception Test mIoU R@0.5 R@0.7 mIoU R@0.5 R@0.7 mIoU acc Overlap R@0.5 Short-Avg Avg Val Baseline Qwen2.5-VL-7B 29.0 24.2 11.1 21.1 15.8 7.5 15.4 59.5 12.6 1.1 71.3 66.9 69.1 SFT on specific tasks +SFT w/ Charades-STA 46.3 45.0 25.3 20.6 16.7 7.9 - - - - N/A* N/A* N/A* +SFT w/ GoT - - - - - - - - 41.8 29.5 59.2 58.6 58.5 +SFT w/ NextQQA - - - - - - 28.2 64.8 - - 60.1 59.2 60.7 GRPO on various tasks VideoChat-R1 60.8 71.7 50.2 36.6 33.4 17.7 32.4 70.6 43.8 38.2 72.2 67.9 70.0 VideoChat-R1-thinking 59.9 70.6 47.2 35.5 33.3 16.7 36.1 69.2 43.3 33.9 74.2 66.2 69.6
@@ 4.2 Ablation Studies and Discussions
Temporal Grounding and Object tracking.
Table 2: Results of Temporal Grounding Task. * indicates that the model has suffered from overfitting and is unable to answer the question properly.
Method GoT VideoMME Average overlap R@0.5 R@0.7 Short-Avg Qwen2.5-VL-7B 12.6 1.1 0 71.3 +SFT 41.8 29.5 3.9 59.2 +GRPO 42.5(+29.2) 30.6(+29.5) 3.9(+3.9) 71.4(+0.1)
Table 3: Results of Object Tracking. We use 8 frames as input for training and evaluation.
Video Question Answer.
Table 4: Results of Multi-Choice Video QA.
Video Caption and Video Quality Assessment.
For the Video Caption and Video Quality Assessment tasks, we found that GRPO still demonstrated its advantages over SFT. Surprisingly, although the reward we designed for the Caption task was not entirely rigorous and reasonable, it still enabled the model to achieve significant improvements in metrics on the Dream-1k benchmark.
Table 5: Results of Video Caption and Video Quality Access.
@@
Table 6: Results of Cotraining on Spatio-Temporal Tasks.
@@ 4.3 Qualitative Results
Figure 3: Examples on Video QA task. It can be seen that VideoChat-R1 can not only answer questions correctly but also provide relatively accurate reference time periods (glue).
As shown in Figure 2 and 3, we provide visualizations of VideoChat-R1’s outputs for the temporal grounding and video QA tasks. Note that VideoChat-R1 possesses strong spatiotemporal perception. Additionally, its thought process exhibits a certain degree of logic. However, due to the limitations of the tasks, compared with the thought processes for math problems in the text and image domains, the thought process of VideoChat-R1 is still rather simplistic and rudimentary.

### v4

@@ 1 Introduction
  • The improvement of spatio-temporal perception ability and the preservation of the original chat capability can contribute to a more reliable and efficient video dialogue system. Our VideoChat-R1 can provide reference video segments when answering users’ questions. Meanwhile, we propose to utilize these video segments for “Clue-Perception" to further obtain more accurate answers. Our experimental results reveal the potential of the approach that enhances the model’s spatiotemporal perception ability through reinforcement learning for future research in the directions of reliable video dialogue systems and long video understanding.
@@ 3.1 Preliminary of Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) [25] is a variant of Proximal Policy Optimization (PPO) [24] in reinforcement learning. By comparing groups of candidates responses directly, GRPO eliminates dependency on a critic model and significantly lowers training resources. Given an input question $q$, GRPO first generates $G$ distinct candidate responses $o=\{o_{1},\dots,o_{G}\}$ through policy sampling. A predefined reward function is used to get the corresponding scores $\{R_{1},\dots,R_{G}\}$. GRPO computes their mean and standard deviation for normalization and determines the quality of these responses:
@@
where $A_{i}$ represents the relative quality of the $i$-th answer. GRPO encourages the model to favor better answers with a high score within the group. The final training objective also considers preventing the optimized policy $\pi_{\theta}$ from deviating far from the original MLLM parameters $\pi_{\mathrm{ref}}$ by adding a KL-divergence term $\mathrm{D}_{\mathrm{KL}}(\cdot|\cdot)$ to :
$\max_{\pi_{\theta}}\mathbb{E}_{o\sim\pi_{\theta_{\text{old}}}}\Big[$ $\sum_{i=1}^{G}\min\Big(r_{i}\cdot A_{i},\ \text{clip}(r_{i},1-\epsilon,1+\epsilon)\cdot A_{i}\Big)$ (2)
$-\ \beta\,\mathrm{D}_{\text{KL}}(\pi_{\theta}\parallel\pi_{\text{ref}})\Big]$
where $r_{i}=\tfrac{\pi_{\theta}(o_{i})}{\pi_{\theta_{\text{old}}}(o_{i})}$, $\beta$ is a regularization coefficient, preventing excessive deviation from the reference policy during optimization.
@@ 3.2 Spatio-Temporal Rewards of Video MLLM in GRPO
$R_{\mathrm{format}}=\begin{cases}1,&\text{if output matches format},\\ 0,&\text{if output doesn't match format}.\end{cases}$ (3)
@@ 3.3 Enhance Spatio-Temporal Perception of Video MLLM through GRPO
VideoChat-R1-18k.
We collect diverse video corpus from existing public video datasets for reinforcement learning training to enhance the model’s spatiotemporal perception ability. For the temporal grounding task, we use the training set of Charade - STA [8] (5,338 samples) for training. For the object tracking task, training is conducted on the GoT - 10k [11] dataset, which has 9,335 samples. For the QA and grounding QA tasks, the validation set of NExTGQA  [35] (3,358 samples) is used for training. For video captioning, FIBER-1k [36] (1,000 samples) is adopted for training. For video quality assessment, we use the quality assessment task from VidTAB [17] under the 100-shot setting, with 200 samples for training. Finally, for the training of VideoChat-R1, we perform joint training on three spatio-temporal perception-related tasks: temporal grounding, object tracking, and grounding QA. In total, 18,031 samples are used for training.
Input: $V_{\text{low}}$: Low-resolution/low-fps video
$Q$: User question
$\Delta_{\text{res}}$: Resolution boost factor
$\Delta_{\text{fps}}$: Frame rate boost factor
Output: $A^{(2)}$: Final refined answer
Initial Answer Generation:
$(A^{(1)},\mathcal{C})\leftarrow\mathcal{M}(V_{\text{low}},Q)$
Clue Processing & Upsampling:
$\mathcal{T}_{c}\leftarrow\text{ExtractTemporalClues}(\mathcal{C})$
$V_{\text{seg}}\leftarrow\text{SelectSegments}(V_{\text{low}},\mathcal{T}_{c})$
$V_{\text{high}}\leftarrow\text{Upsample}(V_{\text{seg}},\Delta_{\text{res}},\Delta_{\text{fps}})$
Final Answer Generation:
$A^{(2)}\leftarrow\mathcal{M}(V_{\text{high}},Q)$
return $A^{(2)}$
Algorithm 1 VideoChat $\mathcal{M}$ with "Clue-Perception"
Chat with enhanced Spatio-Temporal Perception.
After enhancing the spatiotemporal perception capability of MLLMs, we can construct a more reliable video chat system. Specifically, after the model answers a user’s question, it can provide relatively accurate clues that support the answer to that question. We can further leverage these clues to improve the accuracy of the responses. Herein, we propose a simple "Clue-Perception" operation: after the model gives the first answer, we re-input the video segments corresponding to the obtained clues into the model at a higher resolution and frame rate, prompting it to answer again. By perceiving more details, the model can generate more accurate responses. Additionally, this operation is also well-suited for long video understanding under conditions of limited computing resources.
@@ 4.1 Evaluation of VideoChat-R1
Method Charades-STA ActivityNet NExTGQA GoT VideoMME MVBench Peception Test mIoU R@0.5 R@0.7 mIoU R@0.5 R@0.7 mIoU acc Overlap R@0.5 Short-Avg Avg Val Baseline Qwen2.5-VL-7B 29.0 24.2 11.1 21.1 15.8 7.5 15.4 59.5 12.6 1.1 71.3 66.9 69.1 SFT on specific tasks +SFT w/ Charades-STA 46.3 45.0 25.3 20.6 16.7 7.9 - - - - N/A* N/A* N/A* +SFT w/ GoT - - - - - - - - 41.8 29.5 59.2 58.6 58.5 +SFT w/ NExTGQA - - - - - - 28.2 64.8 - - 60.1 59.2 60.7 GRPO on various tasks VideoChat-R1 60.8 71.7 50.2 36.6 33.4 17.7 32.4 70.6 43.8 38.2 72.2 67.9 70.0 VideoChat-R1-thinking 59.9 70.6 47.2 35.5 33.3 16.7 36.1 69.2 43.3 33.9 74.2 66.2 69.6
@@ 4.2 Ablation Studies and Discussions
Table 2: Ablation results of Cotraining on Spatio-Temporal Tasks.
Table 3: Ablation results of Temporal Grounding Task. * indicates that the model has suffered from overfitting and is unable to answer the question properly.
Method GoT VideoMME Average overlap R@0.5 R@0.7 Short-Avg Qwen2.5-VL-7B 12.6 1.1 0 71.3 +SFT 41.8 29.5 3.9 59.2 +GRPO 42.5(+29.9) 30.6(+29.5) 3.9(+3.9) 71.4(+0.1)
Table 4: Ablation results of Object Tracking. We use 8 frames as input for training and evaluation.
Temporal Grounding and Object tracking.
Table 5: Ablation results of Multi-Choice Video QA.
Video Question Answer.
Table 6: Results of Video Caption and Video Quality Access.
Method LLM Judge Dream-1k F1 Precision Recall Baseline - 30.6 33.8 27.9 + GRPO GPT-3.5-turbo-0125 37.9(+7.3) 44.4(+10.6) 33(+5.1) Qwen2.5-72B 38.2(+7.6) 45.4(+11.6) 33.1(+5.2)
Table 7: Ablation of Video Caption Task.
Video Caption and Video Quality Assessment.
For the Video Caption and Video Quality Assessment tasks, we found that GRPO still demonstrated its advantages over SFT, As shown in Table 6. The significant metric improvements on these two benchmarks demonstrate the effectiveness of our approach.
Ablation of Reward Evaluators
To assess the impact of different large language models (LLMs) as reward evaluators, we conducted parallel experiments using GPT-3.5-turbo-0125 and Qwen2.5-72B as distinct judges (Table 7). Models trained under both evaluators achieved nearly identical performance, demonstrating consistent caption reward generation across LLMs. We attribute this consistency to GRPO’s fundamental mechanism: it relies on relative differential scoring within response groups rather than absolute reward values. This confirms that $R_{recall}$ produces discriminative reward signals for predicted captions independent of the choice of LLM judge, validating both the efficacy of our reward design and the stability of its signaling mechanism. Crucially, when guided by these reliable reward signals, our approach delivers substantial performance gains in description tasks using only limited high-quality data, demonstrating remarkable data-efficiency and significant optimization potential.
Model Clue VideoMME LongVideoBench Avg. Duration Perception 1010s 473s Qwen2.5-VL-7B 64.4 56.0 63.3(-1.1) 55.2(-0.8) VideoChat-R1-thinking 62.1 51.9 63.6(+1.5) 58.2(+6.3)
Table 8: Ablation results of "Clue-Perception". It should be noted that due to our adoption of a lower number of input pixels, the absolute performance is not entirely consistent with that reported for Qwen2.5-VL.
Ablation of “Clue-Perception"
As shown in Table. 8, we compared the performance changes of the model with and without perception enhancement when applying the "Clue-Perception" strategy on two representative long video benchmarks [7, 34]. It is noteworthy that without the use of "Clue-Perception", VideoChat-R1 showed no significant performance improvement over Qwen2.5-VL-7B in long video tasks, which can be attributed to the fact that our training dataset consists entirely of short videos under 1 minute. However, after the application of the "Clue-Perception" operation, VideoChat-R1 demonstrated a significant performance enhancement, indicating that the clues it provides are more accurate and thus revealing the potential of clue-perception in long video understanding. In contrast, due to its insufficient spatiotemporal perception capability, Qwen2.5-VL-7B even exhibited a performance decline after the implementation of the "Clue-Perception" operation.
@@
Figure 3: Examples on Video QA task. It can be seen that VideoChat-R1 can not only answer questions correctly but also provide relatively accurate reference time periods (clue).
@@ 4.3 Qualitative Results
As shown in Figure 2 and 3, we provide visualizations of VideoChat-R1’s outputs for the temporal grounding and video QA tasks. We observe that due to the enhanced spatiotemporal perception capability, VideoChat-R1 can more accurately locate video segments relevant to the question and conduct reasonable reasoning to answer the question. However, compared with the complex chains of thought involved in solving mathematical problems or coding tasks in text and image domains, the chains of thought generated in our current video tasks are relatively simple. We will explore more complex video reasoning tasks in future research.